# One-Layer Transformers are Provably Optimal for In-context Reasoning and Distributional Association Learning in Next-Token Prediction Tasks

### v2

@@ Abstract
We study the approximation capabilities and on-convergence behaviors of one-layer transformers on the noiseless and noisy in-context reasoning of next-token prediction. Existing theoretical results focus on understanding the in-context reasoning behaviors for either the first gradient step or when the number of samples is infinite. Furthermore, no convergence rates nor generalization abilities were known. Our work addresses these gaps by showing that there exists a class of one-layer transformers that are provably Bayes-optimal with both linear and ReLU attention. When being trained with gradient descent, we show via a finite-sample analysis that the expected loss of these transformers converges at linear rate to the Bayes risk. Moreover, we prove that the trained models generalize to unseen samples as well as exhibit learning behaviors that were empirically observed in previous works. Our theoretical findings are further supported by extensive empirical validations.
@@ 1 Introduction
Large language models (LLMs) have shown impressive results in complex tasks that require some form of “reasoning” where classical models such as feed-forward networks seem to struggle. These reasoning tasks include, but are not limited to, generating coherent and plausible texts from a given context, language understanding, and mathematical reasoning (Brown et al., 2020; Achiam et al., 2023). At the heart of LLMs is the transformer architecture that features the self-attention mechanism (Vaswani et al., 2017). Transformers can process a long sequence of contexts and enable in-context reasoning via attention mechanisms. Despite such remarkable empirical performance, the theoretical understanding of self-attention in reasoning tasks remains elusive, raising critical risk and safety issues when it comes to the widespread adoption of LLM technology in a socially acceptable manner (Bommasani et al., 2021; Belkin, 2024).
The literature has shown the usefulness of disentangling the behavior of complex models such as LLMs via controlled-setting tasks that we understand the groundtruth behaviors (Allen-Zhu, 2024). For understanding reasoning in LLMs, one of the benchmark tasks that the literature has been recently embarked on is next-token prediction (NTP), wherein the tasks require a model to understand the context from a sentence to be able to predict the next token correcty. As a running example, consider the task of predicting the next token for the sentence “Anna gives a book to” (the same example used in (Chen et al., 2025)). A global bigram statistics would predict the next token to be “the” as the bigram “to the” naturally occurs in English with high frequency. However, if another person appears in the context, say Bob, then “Bob” is perhaps a better token prediction, even though the bigram “to Bob” is not a frequent bigram in the global context. In transformers, there have been strong (both empirically and theoretically) evidence that attention heads seem to responsible for in-context reasoning such as the in-context bigram “to Bob” (Wang et al., 2022) while feedforward layers seem to responsible for storing global statistics or factual knowledge such as the global bigram “to the” (Geva et al., 2020; Meng et al., 2022; Bietti et al., 2023; Nichani et al., 2024).
@@
In this paper, we narrow the gap above by providing a more complete picture of training transformers on next-token prediction tasks and how they get assigned with in-context reasoning and distributional association. In particular, we provide a synthetic next-token prediction task that requires a model to differentiate in-context reasoning from distributional association to be able to succeed. We show that one-layer transformers, from either the class of linear or ReLU attentions, suffice to provably approximate the Bayes-optimal predictor of the task. In addition, we show that training these transformers with normalized gradient descents converges at linear rate to the Bayes optimal predictor and exhibit a provable generalization to unseen samples. Furthermore, we show that, after sufficient number of training iterations, feed-forward layer learns to predict the noise token while attention layers learn to predict the true output token, a behavior that has been observed empirically recently. Finally, we corroborate theoretical findings with extensive empirical validations.
@@ 1.1 Related Work
Brown et al. (2020) demonstrate that trained transformers can perform in-context learning—adapting to new tasks at inference with only a few examples. Building on this, several studies (Akyürek et al., 2022; Bai et al., 2023; Von Oswald et al., 2023; Giannou et al., 2023) show that appropriately parameterized transformers can simulate procedures like gradient descent. Furthermore, (Yun et al., 2019) prove that transformers can approximate any continuous permutation-equivariant sequence-to-sequence function on compact support.
Several works have analyzed transformers’ training dynamics for in-context learning of linear regression. Ahn et al. (2024) show a one-layer linear transformer that performs a preconditioned gradient step, with $L$ layers corresponding to $L$ steps at certain critical points. Mahankali et al. (2023) find that a one-layer linear transformer trained on noisy data mimics a single least-squares gradient step. Zhang et al. (2024) prove convergence to a global minimum under suitable initialization. Huang et al. (2024a) study gradient descent in softmax transformers learning linear functions. Cui et al. (2024) show that multi-head attention with large embeddings outperforms single-head variants. Cheng et al. (2023) demonstrate that nonlinear transformers can emulate gradient descent on nonlinear functions. Kim and Suzuki (2024) analyze mean-field and two-timescale limits of linear attention models, and Chen and Li (2024) give computational bounds for training multi-head attention layers.
Another line of work investigates the training dynamics for binary classification. Tarzanagh et al. (2023) show that self-attention optimization mirrors hard-margin SVMs, revealing the implicit bias of 1-layer transformers trained via gradient descent, and that over-parameterization aids global convergence. Ataee Tarzanagh et al. (2023) demonstrate that gradient descent on softmax attention converges to a max-margin separator distinguishing locally optimal tokens. Building on this, Vasudeva et al. (2024) provide finite-sample analysis. Deora et al. (2023) offer optimization and generalization guarantees for training single-layer multi-head attention models under the NTK regime.
@@ 2 Problem Setup
Definition 2.1.
(Data Model) We study slightly modified variants of the noiseless and noisy in-context reasoning tasks proposed in Bietti et al. (2023) and Chen et al. (2025), respectively. More specifically, we define the following two special, non-overlapping sets of tokens: a set of trigger tokens ${\mathcal{Q}}\subset[N]$ and a set output tokens ${\mathcal{O}}\subset[N]$, where ${\mathcal{O}}\cap{\mathcal{Q}}=\emptyset$. A special “generic” noise token is defined by $\tau=N+1$. The noise level is determined by a constant $\alpha\in[0,1)$, where $\alpha=0$ corresponds to the noiseless learning setting (Bietti et al., 2023) and $\alpha>0$ corresponds to the noisy learning setting (Chen et al., 2025). In our model, a sentence $z_{1:H+1}$ is generated as follows:
  • Sample a trigger word $q\sim\mathrm{Unif}({\mathcal{Q}})$ and an output word $y\sim\mathrm{Unif}({\mathcal{O}})$,
  • Sample randomly (over an arbitrary distribution) $z_{1:H-1}$ from the set of sentences that satisfy the following three conditions: (I) there exists at least one bigram $(q,y)$ in the sentence, (II) $\tau$ may appear in a sentence only if $\alpha>0$, in that case $\tau$ is always preceded by $q$, and (III) no other (trigger, token) bigrams other than $(q,y)$ or $(q,\tau)$ are in the sentence,
@@
Comparisons to existing works. Compared to the existing task modes in Bietti et al. (2023); Chen et al. (2025), our task model offers several notable advantages. First, all sentences in our models must contain at least one (trigger token, output token) bigram, leading to a better signal-to-noise ratio. This allows us to avoid un-informative sentences that contain no useful signals for learning. Second, our task models are agnostic with respect to the distribution of the sentences. In other words, we do not impose any assumptions on how words and sentences are distributed, as long as the conditions are satisfied. Thus, our distributionally agnostic models are both more applicable to practical scenarios and more challenging for theoretical analyses. Third, by restricting the output tokens to a subset of $[N]$, we can study the generalization ability of a model on unseen output tokens. Furthermore, because there are more than one possible next-token for every trigger word, our next-token prediction task is more challenging than the task of learning a one-to-one token mapping in Huang et al. (2024b).
One-layer Decoder-only Transformers. To establish the theoretical guarantees of the optimality and on-convergence behaviors of transformers, we adopt the popular approach in existing works (Bietti et al., 2023; Chen et al., 2025; Huang et al., 2024a, e.g.) and consider the following one-layer transformer model, which is a variant of the model in Chen et al. (2025). Let $E:[N]\mapsto\mathbb{R}^{d}$ be the input word embedding, i.e. $E(z)\in\mathbb{R}^{d}$ is the input embedding of the word $z\in[N]$, and $\tilde{E}:[N]\mapsto\mathbb{R}^{d}$ be a (different) embedding representing the previous token head construction as in Bietti et al. (2023). Let ${\bm{U}}\in\mathbb{R}^{N\times d},{\bm{V}}\in\mathbb{R}^{d\times d},{\bm{W}}\in\mathbb{R}^{d\times d}$ be the unembedding matrix, the value matrix and the joint query-key matrix, respectively. Our model consists of one attention layer and one feed-forward layer. The input ${\mathbf{x}}_{1:H}$ and the output of the model are as below.
@@
$\xi_{F}$ $={\bm{U}}{\bm{F}}\left({\bm{x}}_{H}+\sum_{h=1}^{H}\sigma(x_{H}^{\top}{\bm{W}}{\bm{x}}_{h}){\bm{x}}_{h}\right),$
where $\bm{F}$} is the matrix of the linear layer and $\sigma:\mathbb{R}\to\mathbb{R}$ is the activation function which determines the range of the attention scores. For theoretical analyses, we use linear attention $\sigma({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})={\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h}$ and ReLU attention $\sigma({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})=\max(0,{\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})$. For empirical analyses, we additionally use softmax attention $\sigma({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})=\frac{\exp({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})}{\sum_{j=1}^{H}\exp({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{j})}$. The final logit is $\xi=\xi_{A}+\xi_{F}$.
Compared to Chen et al. (2025), our model in (1) differs in the computation of $\xi_{F}$. More specifically, while their theoretical model used $\xi_{F}={\bm{U}}{\bm{F}}{\bm{x}}_{H}$, we add the input $\sum_{h=1}^{H}\sigma(x_{H}^{\top}{\bm{W}}{\bm{x}}_{h}){\bm{x}}_{h}$ to the feed-forward layer, which is more similar to the empirical model that was used for the experiments in Chen et al. (2025). As we will show in Section 4, this modification is sufficient for showing the Bayes-optimality of one-layer transformers. Next, similar to Chen et al. (2025), we fix the embedding maps $E,\tilde{E},{\bm{U}}$ and use cross-entropy loss on $\xi$, i.e. the population loss is
$L$ $=\mathbb{E}_{q,y,{\bm{z}}}\left[-\ln\frac{\exp(\xi_{y})}{\sum_{j\in[N]}\exp(\xi_{j})}\right].$ (2)
To facilitate the analysis, we employ the following commonly used assumption, which is also the first part of Assumption F.1 in Chen et al. (2025).
Assumption 2.2.
(Orthonormal Embedding) For any $i,j\in[N]$, we have $E(i)^{\top}E(j)=\tilde{E}(i)^{\top}\tilde{E}(j)=\mathbb{1}\{i=j\}$ and $E(i)^{\top}\tilde{E}(j)=0$.
Remark 2.3.
Assumption 2.2 necessarily implies that $d\geq N$. This condition is not too strict as transformers may need sufficient memory capability to remember the vocabulary before they can perform any meaningful in-context reasoning. This is consistent with suggestions in recent literature on the memorization capacity of transformers, where the embedding dimension or the number of heads is polynomially larger than the vocabulary size (Nichani et al., 2024; Mahdavi et al., 2023; Madden et al., 2024).
3 Noiseless Learning ($\alpha=0$)
In this section, we consider the noiseless learning setting in which $\alpha=0$ and $\tau$ never appears in a sentence. In Section˜3.1, we prove the approximation capability of the model defined in (1) by showing that there exists a reparameterization of ${\bm{U}},{\bm{V}},{\bm{F}}$ and $\bm{W}$} that drives the population loss (2) to 0. In Section˜3.2, we show that the reparameterized model can be trained by normalized gradient descent and the population loss converges to 0 at linear rate.
3.1 Approximation Capabilities of One-Layer Transformers On Noiseless Setting
@@
where ${\bm{e}}_{j}$ is the $j$-th vector in the canonical basis of $\mathbb{R}^{N}$ (i.e., ${\bm{e}}_{j}]_{k}=\mathbb{1}\{j=k\}$). Note that the output embedding $\bm{U}$} is considered a fixed matrix as in Chen et al. (2025), thus the population loss is a function of $\bm{V}$} and $\bm{W}$}. We consider a specific parametric class of the weight matrices ${\bm{U}},{\bm{V}}$ and $\bm{W}$}. In particular, the following lemma shows that there exists a reparameterization of ${\bm{U}},{\bm{V}}$ and $\bm{W}$} that makes the population loss arbitrarily close to 0. The proof is in Appendix A.1.
@@
Let $\bm{\lambda}=\{\lambda_{k}\in\mathbb{R}_{+}:k\in{\mathcal{Q}}\}$ be a set of $|\absolutevalue{{\mathcal{Q}}}$ of non-negative values. Under Assumption 2.2, by setting ${\bm{U}}=[E(1)~E(2)~\dots~E(N)]^{\top},{\bm{V}}={\bm{I}}_{d}$ and ${\bm{W}}=\sum_{k\in{\mathcal{Q}}}\lambda_{k}E(k)\tilde{E}^{\top}(k)$, for both linear and ReLU attention, we obtain
$\lim_{\bm{\lambda}\to\infty}L(\bm{\lambda})\vcentcolon=\lim_{(\lambda_{q})_{q\in{\mathcal{Q}}}\to\infty}L\left({\bm{I}}_{d},\sum_{q\in{\mathcal{Q}}}\lambda_{q}E(q)\tilde{E}^{\top}(q)\right)=0.$ (4)
@@
(Sketch) For linear and ReLU attention, it can be shown that the attention score of the $h$-th token $z_{h}$ is $\sigma({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})=\mathbb{1}\{z_{h-1}=q\}$. In other words, the attention scores ${\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h}$ is a non-zero (and positive) value for $x_{h}$ only when $z_{h-1}$ is the trigger word. As a result, the logits of token $j\in[N]$ is $\xi_{j}=\xi_{A,j}=\lambda_{q}C_{q,y}\mathbb{1}\{j=y\}$. It follows that the probability of outputting $y$ is $\lim_{\lambda_{q}\to\infty}\frac{\exp(\xi_{y})}{\sum_{j\in[N]}\exp(\xi_{j})}=\lim_{\lambda_{q}\to\infty}\frac{\exp(C_{q,y}\lambda_{q})}{\exp(C_{q,y}\lambda_{q})+N-1}=1$. ∎
3.2 Training Dynamics and On-Convergence Behavior: A Distributionally Agnostic Analysis
We analyze the gradient descent training dynamics in training one-layer linear transformers parameterized in Section˜3.1. Directly running gradient descent on (3) w.r.t $({\bm{V}},{\bm{W}})$ is difficult due to the fact that the cross-entropy loss may not be jointly-convex with respect to $({\bm{V}},{\bm{W}})$. Previous works (e.g. Ahn et al., 2024; von Oswald et al., 2024) have shown that theoretical insights from the optimization dynamics in linearized transformers generalize well to empirical results of non-linear attentions. Inspired by these works, we focus on characterizing the training dynamics, on-convergence behavior and generalization capability of our model using linear attention. In the following, we show that with linear attention, the population loss converges linearly to zero and the trained model generalizes to unseen output tokens.
Convergence rate of normalized gradient descent. From the proof sketch of Lemma 3.1, the population loss $L(\bm{\lambda})$ is $L(\bm{\lambda})=\mathbb{E}_{q,y,z}\left[-\ln\frac{\exp(C_{q,y}\lambda_{q})}{\exp(C_{q,y}\lambda_{q})+N-1}\right]$. We initialize $\bm{\lambda}_{0}=\mathbf{0}$. Running standard gradient descent $\bm{\lambda}_{t+1}=\bm{\lambda}_{t}-\eta\nabla_{\bm{\lambda}}L$, where $\eta>0$ is the learning rate, would require knowing the exact distribution of $z$ since $C_{q,y}$ is a random variable depending on $z$. Instead, we adopt a normalized gradient descent algorithm, where

### v3

# How Transformers Learn In-Context Recall Tasks? Optimality, Training Dynamics and Generalization

@@ Abstract
We study the approximation capabilities, convergence speeds and on-convergence behaviors of transformers trained on in-context recall tasks – which requires to recognize the positional association between a pair of tokens from in-context examples. Existing theoretical results only focus on the in-context reasoning behavior of transformers after being trained for the one gradient descent step. It remains unclear what is the on-convergence behavior of transformers being trained by gradient descent and how fast the convergence rate is. In addition, the generalization of transformers in one-step in-context reasoning has not been formally investigated. This work addresses these gaps. We first show that a class of transformers with either linear, ReLU or softmax attentions, is provably Bayes-optimal for an in-context recall task. When being trained with gradient descent, we show via a finite-sample analysis that the expected loss converges at linear rate to the Bayes risks. Moreover, we show that the trained transformers exhibit out-of-distribution (OOD) generalization, i.e., generalizing to samples outside of the population distribution. Our theoretical findings are further supported by extensive empirical validations, showing that without proper parameterization, models with larger expressive power surprisingly fail to generalize OOD after being trained by gradient descent.
@@ 1 Introduction
Large language models (LLMs) have shown impressive results in complex tasks that require some form of “reasoning” where classical models such as feed-forward networks seem to struggle. These reasoning tasks include, but are not limited to, generating coherent and plausible texts from a given context, language understanding, and mathematical reasoning (Brown et al., 2020; Achiam et al., 2023). At the heart of LLMs is the transformer architecture that features the attention mechanism (Vaswani et al., 2017). Transformers can process a long sequence of contexts and enable in-context reasoning via attention mechanisms. Despite remarkable empirical performance, the theoretical understanding of attention in reasoning tasks remains elusive, raising critical risk and safety issues when it comes to the widespread adoption of LLM technology (Bommasani et al., 2021; Belkin, 2024).
The literature has shown the usefulness of disentangling the behavior of complex models such as LLMs via controlled-setting tasks that we understand the groundtruth behaviors (Allen-Zhu, 2024). For understanding reasoning in LLMs, one of the benchmark tasks that the literature has been recently embarked on is next-token prediction (NTP), wherein the tasks require a model to understand the context from a sentence to be able to predict the next token correcty. As a running example, consider the task of predicting the next token for the sentence “After talking to Bob about Anna, Charles gives her email address to [?]”. A global bigram statistics would predict the next token to be “the” as the bigram “to the” naturally occurs in English with high frequency. However, if another person appears in the context, say Bob, then “Bob” is perhaps a better token prediction, even though the bigram “to Bob” is not a frequent bigram in the global context. In transformers, there have been strong (both empirically and theoretically) evidence that attention heads seem to responsible for in-context reasoning such as the in-context bigram “to Bob” (Wang et al., 2022) while feedforward layers seem to responsible for storing global statistics or factual knowledge such as the global bigram “to the” (Geva et al., 2020; Meng et al., 2022; Bietti et al., 2023; Nichani et al., 2024).
@@
In this paper, we narrow the gap above by deriving an effective, interpretable structures of transformers for in-context recall tasks. We will show how these structures emerge through gradient descent training on a class of parameterized one-layer transformers with linear, ReLU and softmax attentions. Our main contributions are as follows.
  • In Section 2, we formally define a new in-context reasoning task (Definition 2.1), in which multiple query tokens can appear in a sentence and the output tokens can be noisy. This is a more difficult version of the in-context recall tasks in existing works (Bietti et al., 2023; Chen et al., 2025). Our new data model also enables a natural way to setup out-of-distribution (OOD) testing via a set of neutral tokens.
  • Section 3 considers the noiseless setting. In Lemma 3.1, we present the first parameterization of one-layer transformers with linear and ReLU attentions that are provably optimal for this setting. We show that this parameterization mimics a human-like strategy for solving the task, and that it can be realized via gradient descent training (Theorem 3.2). Furthermore, we prove that the trained model directly generalizes to OOD sentences (Theorem 3.3), as well as gradient descent alone is not implicitly biased towards this parameterization (Theorem 3.4).
  • Section 4 studies the optimality and convergence of models with softmax attentions for the noiseless setting. Lemma 4.1 exhibits how the structure for softmax attention can be constructed from observing the structure for linear and ReLU attentions. Via a two-phase analysis, our Theorem 4.2 shows that the loss converges at a linear rate to 0.
  • Section 5 studies the noisy setting. Lemmas 5.1 and 5.2 first demonstrate the Bayes-optimality of one-layer transformers with linear, ReLU and softmax attentions. Next, Theorem 5.4 shows that by adapting our parameterizations from the noiseless setting to the noisy setting, one-layer transformers admit a finite-sample analysis that results in a PAC-style high-probability generalization bound. Moreover, Theorem 5.6 explains how attention layer and feed-forward layer may converge to perform different functionalities.
  • Finally, Section 6 presents experimental results demonstrating the advantages of our parameterization over non-parameterized models. These results reveal the crucial role of expressive power and parameterization in achieving Bayes-optimality and OOD generalization.
@@ 1.1 Related Work
Several works have analyzed transformers’ training dynamics for in-context learning of linear regression and binary classification. Ahn et al. (2024) show a one-layer linear transformer that performs a preconditioned gradient step, with $L$ layers corresponding to $L$ steps at certain critical points. Mahankali et al. (2023) find that a one-layer linear transformer trained on noisy data mimics a single least-squares gradient step. Zhang et al. (2024) prove convergence to a global minimum under suitable initialization. Huang et al. (2024a) study gradient descent in softmax transformers learning linear functions. Cui et al. (2024) show that multi-head attention with large embeddings outperforms single-head variants. Cheng et al. (2023) demonstrate that nonlinear transformers can emulate gradient descent on nonlinear functions. Siyu et al. (2024) studies the training dynamics of multi-head softmax transformers for multi-task linear regression. Tarzanagh et al. (2023) show that self-attention optimization mirrors hard-margin SVMs, revealing the implicit bias of 1-layer transformers trained via gradient descent, and that over-parameterization aids global convergence. Ataee Tarzanagh et al. (2023) demonstrate that gradient descent on softmax attention converges to a max-margin separator distinguishing locally optimal tokens. Building on this, Vasudeva et al. (2024) provide finite-sample analysis. Deora et al. (2023) offer optimization and generalization guarantees for training single-layer multi-head attention models under the NTK regime.
@@ 2 Problem Setup
Definition 2.1 (Data Model - In-context Recall Tasks).
We study slightly modified variants of the noiseless and noisy in-context reasoning tasks proposed in Bietti et al. (2023) and Chen et al. (2025), respectively. More specifically, we define the following two special, non-overlapping sets of tokens: a set of trigger tokens ${\mathcal{Q}}\subset[N]$ and a set output tokens ${\mathcal{O}}\subset[N]$, where ${\mathcal{O}}\cap{\mathcal{Q}}=\emptyset$. A special “generic” noise token is defined by $\tau=N+1$. The noise level is determined by a constant $\alpha\in[0,1)$, where $\alpha=0$ corresponds to the noiseless learning setting (Bietti et al., 2023) and $\alpha>0$ corresponds to the noisy learning setting (Chen et al., 2025). In our model, a sentence $z_{1:H+1}$ is generated as follows:
  • Sample a trigger word $q\sim\mathrm{Unif}({\mathcal{Q}})$ and an output word $y\sim\mathrm{Unif}({\mathcal{O}})$.
  • Sample randomly (over an arbitrary distribution) $z_{1:H-1}$ from the set of sentences that satisfy the following four conditions: (I) there exists at least one bigram $(q,y)$ in the sentence, (II) $\tau$ may appear in a sentence only if $\alpha>0$, in that case $\tau$ is always preceded by $q$, (III) all bigrams of the form $(q,x)$ take either $x=y$ or $x=\tau$, and (IV) if another token $q^{\prime}$ is in the sentence, then it is followed by an output word $y^{\prime}\in{\mathcal{O}}$.
@@
Comparisons to existing works. Compared to the existing task modes in Bietti et al. (2023); Chen et al. (2025), our task model offers several notable advantages. First, all sentences in our models must contain at least one (trigger token, output token) bigram, leading to a better signal-to-noise ratio. This allows us to avoid un-informative sentences that contain no useful signals for learning. Second, our task models are agnostic with respect to the distribution of the sentences. In other words, we do not impose any assumptions on how words and sentences are distributed, as long as the conditions are satisfied. Thus, our distributionally agnostic models are both more applicable to practical scenarios and more challenging for theoretical analyses. Third, by restricting the output tokens to a subset of $[N]$, we can study the OOD generalization ability of a model on unseen output tokens. Furthermore, because there are more than one possible next-token for every trigger word, our next-token prediction task is more challenging than the task of learning a one-to-one token mapping in Huang et al. (2024b). We provide additional examples of real-life sentences in Appendix A.
One-layer Decoder-only Transformers. To establish the theoretical guarantees of the optimality and on-convergence behaviors of transformers, we adopt the popular approach in existing works (Bietti et al., 2023; Chen et al., 2025; Huang et al., 2024a, e.g.) and consider the following one-layer transformer model, which is a variant of the model in Chen et al. (2025). Let $E:[N]\mapsto\mathbb{R}^{d}$ be the input word embedding, i.e. $E(z)\in\mathbb{R}^{d}$ is the input embedding of the word $z\in[N]$, and $\tilde{E}:[N]\mapsto\mathbb{R}^{d}$ be a (different) embedding representing the previous token head construction as in Bietti et al. (2023). Similar to the majority of existing works in the literature (Chen et al., 2025, e.g.), we employ a common assumption that the embeddings $E$ and $\tilde{E}$ are fixed and orthogonal, i.e. $E(i)^{\top}E(j)=\tilde{E}(i)^{\top}\tilde{E}(j)=\mathbb{1}\{i=j\}$ and $E(i)^{\top}\tilde{E}(j)=0$ for any $i,j\in[N]$.
Let ${\bm{U}}\in\mathbb{R}^{N\times d},{\bm{V}}\in\mathbb{R}^{d\times d},{\bm{W}}\in\mathbb{R}^{d\times d}$ be the unembedding matrix, the value matrix and the joint query-key matrix, respectively. Our model consists of one attention layer and one feed-forward layer. The input ${\mathbf{x}}_{1:H}$ and the output of the model are as below.
@@
$\xi_{F}$ $={\bm{U}}{\bm{F}}\left({\bm{x}}_{H}+\phi({\bm{x}}_{H},{\bm{x}}_{1:H})\right)\in\mathbb{R}^{N},$
where $\bm{F}$} is the matrix of the linear layer and $\sigma:\mathbb{R}\to\mathbb{R}$ is the activation function which determines the range of the attention scores. For theoretical and empirical analyses, we use linear attention $\sigma({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})={\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h}$, ReLU attention $\sigma({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})=\max(0,{\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})$ and softmax attention $\sigma({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})=\frac{\exp({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})}{\sum_{j=1}^{H}\exp({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{j})}$. The final logit is $\xi=\xi_{A}+\xi_{F}$.
Compared to Chen et al. (2025), our model in (1) differs in the computation of $\xi_{F}$. More specifically, while their theoretical model used $\xi_{F}={\bm{U}}{\bm{F}}{\bm{x}}_{H}$, we add $\sum_{h=1}^{H}\sigma(x_{H}^{\top}{\bm{W}}{\bm{x}}_{h}){\bm{V}}{\bm{x}}_{h}$ to the input of the feed-forward layer, which is closer to the empirical model that was used for the experiments in Chen et al. (2025). As we will show in Section 5, this modification is sufficient for showing the Bayes-optimality of one-layer transformers. Next, similar to Chen et al. (2025), we fix the embedding maps $E,\tilde{E},{\bm{U}}$ and use cross-entropy loss on $\xi$, i.e. the population loss is $L=\mathbb{E}_{q,y,{\bm{z}}}\left[-\ln\frac{\exp(\xi_{y})}{\sum_{j\in[N]}\exp(\xi_{j})}\right]$.
3 Warmup: Noiseless Learning with Linear and ReLU attentions
In this section, we consider the noiseless learning setting in which $\alpha=0$ and $\tau$ never appears in a sentence. In Section˜3.1, we prove the approximation capability of the model defined in (1) by showing that with linear and ReLU attentions, there exists a reparameterization of ${\bm{U}},{\bm{V}},{\bm{F}}$ and $\bm{W}$} that drives the population loss $L$ to 0. In Section˜3.2, we show that the reparameterized model can be trained by normalized gradient descent (NGD) and the population loss converges to 0 at linear rate.
3.1 Approximation Capabilities of Transformers On Noiseless Setting
@@
where ${\bm{e}}_{j}$ is the $j$-th vector in the canonical basis of $\mathbb{R}^{N}$ (i.e., ${\bm{e}}_{j}]_{k}=\mathbb{1}\{j=k\}$). Note that the output embedding $\bm{U}$} is considered a fixed matrix as in Chen et al. (2025), thus the population loss is a function of $\bm{V}$} and $\bm{W}$}. We consider a specific parametric class of the weight matrices ${\bm{U}},{\bm{V}}$ and $\bm{W}$}. In particular, the following lemma shows that there exists a reparameterization of ${\bm{U}},{\bm{V}}$ and $\bm{W}$} that makes the population loss arbitrarily close to 0. The proof is in Appendix B.1.
@@
Let $\bm{\lambda}=\{\lambda_{k}\in\mathbb{R}_{+}:k\in{\mathcal{Q}}\}$ be a set of $|\absolutevalue{{\mathcal{Q}}}$ of non-negative values. By setting ${\bm{U}}=[E(1)~E(2)~\dots~E(N)]^{\top},{\bm{V}}={\bm{I}}_{d}$ and ${\bm{W}}=\sum_{k\in{\mathcal{Q}}}\lambda_{k}E(k)\tilde{E}^{\top}(k)$, for both linear and ReLU attention, we obtain $\lim_{\bm{\lambda}\to\infty}L(\bm{\lambda})\vcentcolon=\lim_{(\lambda_{q})_{q\in{\mathcal{Q}}}\to\infty}L\left({\bm{V}},{\bm{W}}\right)=0.$
@@
(Sketch) For linear and ReLU attention, it can be shown that the attention score of the $h$-th token $z_{h}$ is $\sigma({\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h})=\lambda_{q}\mathbb{1}\{z_{h-1}=q\}$. In other words, the attention scores ${\bm{x}}_{H}^{\top}{\bm{W}}{\bm{x}}_{h}$ is a non-zero (and positive) value for $x_{h}$ only when $z_{h-1}$ is the trigger word. As a result, the logits of token $j\in[N]$ is $\xi_{j}=\xi_{A,j}=\lambda_{q}C_{q,y}\mathbb{1}\{j=y\}$. It follows that the probability of outputting $y$ is $\lim_{\lambda_{q}\to\infty}\frac{\exp(\xi_{y})}{\sum_{j\in[N]}\exp(\xi_{j})}=\lim_{\lambda_{q}\to\infty}\frac{\exp(C_{q,y}\lambda_{q})}{\exp(C_{q,y}\lambda_{q})+N-1}=1$. ∎
3.2 Convergence rate, Generalization and Implicit Bias of Gradient Descent
We analyze the dynamics of normalized gradient descent (NGD) in training one-layer transformers parameterized in Section˜3.1. We will show that with linear and ReLU attentions, the population loss converges linearly to zero. Moreover, the trained model generalizes to samples that lie completely outside of the training population.
Convergence rate of NGD. From the proof sketch of Lemma 3.1, the population loss $L(\bm{\lambda})$ is $L(\bm{\lambda})=\mathbb{E}_{q,y,z}\left[-\ln\frac{\exp(C_{q,y}\lambda_{q})}{\exp(C_{q,y}\lambda_{q})+N-1}\right]$. We initialize $\bm{\lambda}_{0}=\mathbf{0}$. Running standard gradient descent $\bm{\lambda}_{t+1}=\bm{\lambda}_{t}-\eta\nabla_{\bm{\lambda}}L$, where $\eta>0$ is the learning rate, would require knowing the exact distribution of $z$ since $C_{q,y}$ is a random variable depending on $z$. Instead, we adopt a NGD algorithm, where